An average of 10 cars per hour arrive to the bank at a single server drive-in teller. Assume that the average service time for each customer is 4 minutes, and both interarrival and service times are exponential. What is the probability that the arriving car waits for the teller. 



1
Expert's answer
2021-07-06T12:06:04-0400

If the arriving car waits for the teller, It means that the teller is working rather than being idle.

"M\/M\/1" with "\lambda=10cars/hour" and "\mu =15cars/hour"


Therefore, the probability that the arriving car waits the teller is

"P\left(car\:waits\:for\:teller\right)=\frac{\lambda }{\mu }=\frac{10}{15}"


"P\left(car\:waits\:for\:teller\right)=\frac{2}{3}"